OpenStudy (anonymous):

11*11=4 22*22=16 33*33=??

OpenStudy (anonymous):

18. It is the sum of digits in square of the given number(Here 33). 33*33=1089 1+8+9=18

OpenStudy (anonymous):

33*33= 81 [(3*3) * (3*3) = 81] :)